OpenStudy (anonymous):

A hockey puck is hit on a frozen lake and starts moving with a velocity of 12.0 m/s. 5.0 seconds later, the velocity of the puck is 6.0 m/s. What is the coefficient of friction between the puck and the ice?

OpenStudy (anonymous):

ANY MULTIPLE CHOICE

OpenStudy (anonymous):

@riojabro

OpenStudy (shamim):

Can u tell me the retardation of ur puck

OpenStudy (michele_laino):

we can apply this formula: \[\Large R\Delta t = m\Delta v\;\;\;\;\;\;\;\; (1)\] where \(R\) is the magnitude of the friction force. So, we can write: \[\Large R = \mu mg \;\;\;\;\;\;\;\; (2)\] where \(\mu\) is the friction coefficient. Combining equation \(2\) with equation \(1\), we get: \[\Large \mu mg = \frac{{m\Delta v}}{{\Delta t}}\] therefore: \[\Large \mu = \frac{{\Delta v}}{{g\Delta t}} = \frac{{12 - 6}}{{9.81 \cdot 5}} = ...?\]

OpenStudy (gebooors):

Also friction force is opposite to velocity and kinetic energy decreases. µ has no unit.